Can someone please help with the following, and explain formula to get answer, I'm so lost.
Circumference = 12 meters, length of arc AB = 4 meters, find the measure of the central angle ACB.
and
circumference = 18 feet, measure of angle ACB = 60 degrees. Find the length of arc BD.

Answer by Alan3354(69443)

Circumference = 12 meters, length of arc AB = 4 meters, find the measure of the central angle ACB.
-------
The arc = 1/3 of the circumference, so the angle = 1/3 of 360 = 120 degrees
---------------
and
circumference = 18 feet, measure of angle ACB = 60 degrees. Find the length of arc BD.
-----------
Arc = 18*(60/360) = 3 feet